OpenStudy (anonymous):

For their class, a graduate professor asks students to study an average of 3 hours a day every weekday. If a particular student studied for 1 hour on Monday and 3 hours on Tuesday, how many more hours should they study throughout the week to achieve the desired average?

OpenStudy (anonymous):

11 hours?

OpenStudy (anonymous):

I did that wrong. 11 hours. XD 1 +3 = 4. 3 * 5 = 15 15 - 4 = 11

OpenStudy (anonymous):

its ok want to do another?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ill post again so you get some medals

OpenStudy (anonymous):

I'm not awake..it's like 3 : 15 AM . Sure.

OpenStudy (anonymous):

cool